OpenStudy (anonymous):

evaluate 12∑n=3 -5n-1 help me step by step please

OpenStudy (anonymous):

did u try typing the exact same thing on google...usually answer.com give you good answers :)

OpenStudy (anonymous):

hold on..let me see though

OpenStudy (anonymous):

actually have, but nothing came up for me :(

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

something is not clear in the question

OpenStudy (misty1212):

\[12\sum_{n=3}^?(-5n-1)\]

OpenStudy (anonymous):

the 12 is where the question mark is

OpenStudy (misty1212):

ooh

OpenStudy (misty1212):

\[\sum_{n=3}^{12}(-5n-1)\]

OpenStudy (anonymous):

exactly that lol. can you explain it to me?

OpenStudy (misty1212):

you have to add up a bunch of stuff

OpenStudy (anonymous):

guide me through it step by step?

OpenStudy (misty1212):

we can break it apart in to a couple pieces

OpenStudy (anonymous):

alright :)

OpenStudy (misty1212):

\[-5\sum_{n=3}^{12}n-\sum_{n=3}^{12}1\]

OpenStudy (anonymous):

im confused, what method is this? this wasnt in the lesson :o

OpenStudy (misty1212):

ok then lets do it a different way

OpenStudy (misty1212):

the first number is what you get when \(n=3\)

OpenStudy (misty1212):

\[-5\times 3-1=-16\]

OpenStudy (misty1212):

the last number is what you get when \(n=12\) \[-3\times 12-1=-37\]

OpenStudy (misty1212):

there are 10 numbers between 3 and 12 inclusive

OpenStudy (misty1212):

so you can use \[\frac{10}{2}\left(-16-37)\right)\]

OpenStudy (misty1212):

i messed up womewhere hold on

OpenStudy (anonymous):

if i use that formula, will it give me the total of the sum?

OpenStudy (misty1212):

the last number is what you get when \(n=12\) should be \[-5\times 12-1=-61\]

OpenStudy (misty1212):

that looks better

OpenStudy (misty1212):

\[\frac{10}{2}\left(-16-61\right)\]

OpenStudy (misty1212):

http://www.wolframalpha.com/input/?i=5%28-16-61%29

OpenStudy (misty1212):

here it is without the formula, same answer from wolfram http://www.wolframalpha.com/input/?i=sum+n+%3D+3+to+12%28-5n-1%29

OpenStudy (anonymous):

ohhhhhh, thanks a lot thats the answer i got. can you help me another one?

OpenStudy (misty1212):

k

OpenStudy (anonymous):

5∑ n=1 3(-2)^n-1 i got 33 but i wanna check and make sure if i did this correctly

OpenStudy (misty1212):

is the five on top?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

33 is right

OpenStudy (anonymous):

how? can i see the steps?

OpenStudy (misty1212):

however you did it

OpenStudy (misty1212):

5 is a small enough number you can just add

OpenStudy (anonymous):

alright thanks, can you help with one last question?

OpenStudy (misty1212):

\[3\times 1+3\times (-2)+3\times 4+3\times (-8)+3\times 16\]

OpenStudy (misty1212):

ok but this one will cost you

OpenStudy (misty1212):

jk

OpenStudy (anonymous):

LOL thanks a lot really!~ i have tons of doubts on this one really. given the sequence 5,-10,20... determine the sigma notation of the sum for term 4 through 15

OpenStudy (anonymous):

its asking for the formula on what would it be

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

\[\sum_{n=4}^{15}5\times (2)^{n-1}\] should do it

OpenStudy (misty1212):

on no

OpenStudy (misty1212):

\[\huge \sum_{n=4}^{15}5\times (2)^{n-4}\]

OpenStudy (misty1212):

check to see if the first three terms are right

OpenStudy (anonymous):

that isnt among the options they give me for this question. the 2 in the parenthesis is negative and the exponent is n-1

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

@misty1212 are you still there?

OpenStudy (anonymous):

@mathrulezz

OpenStudy (anonymous):

im sorry! i dont' know how to do those :(